OpenStudy (anonymous):

Parallel and Perpendicular Lines: 1. Write the equation of the line which passes though (-4, 2) and is parallel to y = -x + 6 in slop-intercept form. 2. Write the equation of the line which passes though (2, -3) and is perpendicular to y = 4x + 7 in slope-intercept form.

OpenStudy (anonymous):

@MAYSTEPH

OpenStudy (anonymous):

1. y=-x-2

OpenStudy (anonymous):

How did you get that?

OpenStudy (anonymous):

I got this. . .

OpenStudy (anonymous):

y = mx + b y = -x + 6 (Slope = m = 0) y – 2 = 0(x - -4) y – 2 = 0x + 0 y = 2

OpenStudy (anonymous):

you used the worng equation. I took the original problem to get the slope.

OpenStudy (anonymous):

if you use the equation y-y1=m(x-x1) you also get the original problem and you cant have a slope of zero.

OpenStudy (anonymous):

and since the lines to be parrallel you have to have the same slope for both equations and i new that a negitive times a negitive isa positive i was able to find out that y=-(-4) would give me y=4 and knowing that the 4 subtracted by 2 would leave me with y=2 I was able to figure out that the equation of a parralle line is y=-x-2

OpenStudy (anonymous):

for the second one I know to have a perpendicular line I have to take the original equation and switch the slope from 4 to -1/4 so the I can use the equation y-y1=m(x-x1) so y-(-3)=-1/4(x-2) y+3=-1/4x+1/2 subtract 3 from both sides

OpenStudy (anonymous):

2. y=-1/4x-5/2

OpenStudy (anonymous):

1. y=-x-2 2. y=-1/4x-5/2